OpenStudy (anonymous):

Zach ran t times around the track at school, which is 1/4 mile. He ran a total of 7/8 mile. Solve for t. 3 1/2 7/32 2/7 4 4/7

OpenStudy (anonymous):

(1/4) t = 7/8 t = 7*4 / 8 = 28/8 = 3 1/2